An airplane propeller rotates 1000 times per minute. Find the number of degrees that a point on the edge of the propeller will rotate in 1 second.

Found 2 solutions by Theo, stanbon:
Answer by Theo(13342)

The propeller rotate4s 1000 times per minute.

That means that the propeller rotates 1000 / 60 times per second.

That's 16 and 2/3 revolutions in 1 second.

16 * 360 = 5760 degrees.

2/3 * 360 = 240 degrees.

Total number of degrees is equal to 5760 + 240 = 6000 degrees per second.

You could also have taken 1000 * 360 = 360000 degrees per minutes which, when divided by 60, equals 6000 degrees per second.

1 revolution is 360 degrees.
2 revolutions is 2 * 360 = 720 degrees.
3 revolutions is 3 * 360 = 1080 degrees.
.....
1000 revolutions is 1000 * 360 = 36000 degrees.
etc.....




Answer by stanbon(75887)

An airplane propeller rotates 1000 times per minute. Find the number of degrees that a point on the edge of the propeller will rotate in 1 second.
--------------------
(1000*360)(degrees/minute)(1/3600)(minute/seconds)
---
= (360000/3600)(degrees/second)
---
= 100 degrees/second
==========================
Cheers,
Stan H.